On a vacation, Johny averaged 50 mph traveling from Denver to Minneapolis. Returning by a different route that covered the same number of miles, he averaged 50 mph. What is the distance between the two cities if his total traveling time is 32 hours.
Answer by ankor@dixie-net.com(22740)

On a vacation, Johny averaged 50 mph traveling from Denver to Minneapolis. Returning by a different route that covered the same number of miles, he averaged 50 mph. What is the distance between the two cities if his total traveling time is 32 hours.
:
Let d = distance between the two cities:
:
write a time equation; time = distance/speed
:
Den/Minn time + Minn/Den = 32 hrs
d/55 + d/50 = 32
:
Multiply equation by 550 to get rid of the denominators, you then have:
10d + 11d = 550(32)
21d = 17600
d = 17600/21
d = 838.1 miles
:
:
Check solution of d = 838.1
838.1/55 + 838.1/50 =
15.24 + 16.76 = 32 hrs